OpenStudy (anonymous):

The width of a rectangle is 0.6 of the length. If the width is 9 ft, what is the length? I desperately need help with the question. Please give a helpful and thorough explanation and answer.

OpenStudy (anonymous):

This is the information you're given written algebraically: W=width L= length W=0.6 * L W=9 Therefore, you can figure out that 9 = 0.6 L or L=9/0.6 or L=15

OpenStudy (anonymous):

Oh, okay, thanks! So, what is the algebraic equation?

OpenStudy (snowfire):

.6 means it is 60% of the length. Just multiply the length by .6 or 60%, whichever you are more comfortable with, and that is your width: .6*L=W Just do some algebra and you get: L=W/(.6)